# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 3.1.3 Continuous Dynamics Modeling
RNA Velocity Model and Parameter Estimation
Based on the unspliced RNA and spliced RNA counts $u_{g}$ and $s_{g}$ for each gene, an underlying ODE model could be naturally derived based on mass-action law such that
$\dfrac{\mathrm{d}u_{g}}{\mathrm{d}t}$ $=\alpha_{g}(t)-\beta_{g}(t)u_{g}(t),$ (3.4)
@@
where $\alpha_{g}(t)$, $\beta_{g}(t)$, and $\gamma_{g}(t)$ represent the rates of mRNA transcription, splicing, and degradation respectively. Here, $v_{g}=\dfrac{\mathrm{d}s_{g}}{\mathrm{d}t}$ is defined as the RNA velocity of gene $g$ (La Manno et al., 2018). By concatenating the RNA velocities of all genes in a cell, a vector $\mathbf{v}=(v_{1},v_{2},\cdots,v_{n})$ is formed, which contains information about how the amounts of spliced RNA in the cell are changing. This vector represents the potential direction of the cell state evolution and can be used for downstream tasks of cell fate inference.
From an algorithm perspective, the central issue in the RNA velocity is to determine the parameters of the equation (3.4) from static snapshot data, where the time $t$ of each cell is not explicitly known. In the following sections, we will summarize methods for solving $v_{g}$ and downstream tasks that utilize $v_{g}$.
Steady-state Assumption: Parameter Estimation in Velocyto (La Manno et al., 2018)
In the original RNA velocity paper (La Manno et al., 2018), parameter estimation was performed using linear regression with steady-state assumption to avoid the reliance on latent time $t$. Firstly, it assumes that for all genes $g$, $\alpha_{g}(t)$, $\beta_{g}(t)$, and $\gamma_{g}(t)$ are time-invariant. Secondly, it assumes that all genes share the same splicing rate $\beta$. Denote $\tilde{\alpha}=\dfrac{\alpha}{\beta}$ and $\tilde{\gamma}=\dfrac{\gamma}{\beta}$. To compute RNA velocity, one only needs to estimate $\tilde{\gamma}_{g}$ with the steady state assumption that $\frac{ds_{g}}{dt}=0$. Indeed we have the linear relation $\tilde{\gamma}_{g}=\dfrac{u_{g}(t)}{s_{g}(t)},$suggesting that under the steady-state assumption, $\tilde{\gamma}_{g}$ can be estimated using linear regression. In practice, since most cells do not satisfy this assumption, it is commonly assumed that cells in the upper-right or lower-left regions of a scatter plot with unspliced RNA on the x-axis and spliced RNA on the y-axis are in equilibrium. Therefore, the common algorithm implementation is to limit the linear regression to the top or bottom $5\%$ of cells based on unspliced and spliced RNA levels.
If the dynamic equations Eq. 3.4 are expressed probabilistically, the parameter estimation could be enhanced based on the linear regression formulation of steady-state stochastic model (Bergen et al., 2020). The equation could be expressed as the regression problem
$\begin{bmatrix}\langle u_{g}(t)\rangle\\ \langle u_{g}(t)\rangle+2\langle u_{g}(t)s_{g}(t)\rangle\end{bmatrix}=\tilde{\gamma}_{g}\begin{bmatrix}\langle s_{g}(t)\rangle\\ 2\langle s_{g}^{2}(t)\rangle-\langle s_{g}(t)\rangle\end{bmatrix}+\mathbf{\epsilon},$
@@
Dynamic Inference: Parameter Estimation in scVelo (Bergen et al., 2020)
A major issue with steady-state analysis is that many transient cells would be discarded in the parameter estimation. The scVelo approach (Bergen et al., 2020) circumvents the issues through the estimation of kinetic parameters ($\alpha_{g}$, $\beta_{g}$, $\gamma_{g}$) via an expectation-maximization (EM) algorithm by modeling all the cells with the dynamical process.
Guided by transcriptional regulation principles, scVelo models gene expression dynamics through two distinct transcriptional phases: 1) an induction phase ($k=0$) characterized by promoter activation and transcriptional upregulation, and 2) a subsequent repression phase ($k=1$) marked by transcriptional suppression. This phase-specific regulation manifests through different unspliced RNA production rates ($\alpha^{(0)}_{g}\neq\alpha^{(1)}_{g}$). Let $t^{(k)}_{g}$ denote the transition time from phase $k-1$ to $k$ for gene $g$, with initial conditions $u_{g}^{\star}=u_{g}(t^{(k)}_{g})$ and $s_{g}^{\star}=s_{g}(t^{(k)}_{g})$. The analytical solution to Eq. 3.4 during phase $k$ yields
@@
For each gene $g$, one can estimate the parameter set $\mathbf{\theta}_{g}=\{\alpha^{(k)}_{g},\beta_{g},\gamma_{g},t^{(k)}_{g}\}$ by minimizing the discrepancy between modeled trajectories $\mathbf{\hat{x}}_{g}(t)=(u_{g}(t),s_{g}(t))$ and observed single-cell measurements $\mathbf{x}_{g,c}=(u_{g,c},s_{g,c})$ across cells $c$. Assuming Gaussian residuals $\mathbf{e}_{c}=\|\mathbf{x}_{g,c}-\mathbf{\hat{x}}_{g}(t_{c})\|$ with variance $\sigma^{2}$, the log-likelihood function becomes
$\max_{\mathbf{\theta_{g}},t_{c}}\mathcal{L}(\mathbf{\theta}_{g})=-\frac{1}{2\sigma^{2}}\sum_{c}\|\mathbf{x}_{g,c}-\mathbf{\hat{x}}_{g}(t_{c})\|^{2}+\text{constant}.$ (3.5)
@@
    $\beta_{g}$ $=1,\quad\gamma_{g}=\frac{\mathbf{u}_{g}^{\top}\mathbf{s}_{g}}{\|\mathbf{s}_{g}\|^{2}},$
@@
  • E-step: Assigning hidden latent time $t_{c}$ for each cell by projecting observations onto the current estimated trajectory $\mathbf{\hat{x}}_{g}(t|\mathbf{\theta}_{g})$.
  • M-step: Updating $\mathbf{\theta}_{g}$ via maximum likelihood estimation given current latent time assignments.
@@
While traditional RNA velocity methods focus on estimating parameters $\alpha_{g},\beta_{g},\gamma_{g}$ in dynamic equations (3.4), alternative approaches such as UniTVelo (Gao et al., 2022) and TF Velo (Li et al., 2024a) take a different path by directly parameterizing the dynamics of spliced RNA. UniTVelo (Gao et al., 2022) models transcriptional phases through radial basis functions
@@
where the velocity derives directly from function differentiation $\dot{s}_{g}(t_{g,c})=-2a_{g}(t_{g,c}-\tau_{g})s_{g}(t_{g,c}).$ The full parameter set $(h_{g},a_{g},\tau_{g},o_{g},\gamma_{g},\beta_{g},i_{g},t_{g,c})$ is estimated via maximum likelihood framework as Equation (3.5), comparing model predictions $\mathbf{\hat{x}}_{c}=(u_{g}(t_{g,c}),s_{g}(t_{g,c}))$ against observations $\mathbf{x}_{c}=(u_{g,c},s_{g,c})$ under Gaussian residuals.
TF Velo (Li et al., 2024a) introduces transcription factor coupling through linear dynamics $\dot{s}_{g}(t)=\mathbf{w}_{g}^{\top}\mathbf{f}_{g}(t)-\gamma_{g}s_{g}(t),$ combined with sinusoidal splicing $s_{g}(t)=A_{g}\sin(\omega_{g}t+\theta_{g})+b_{g}.$ This functional specification enables the analytical resolution of TF interactions
$\mathbf{w}_{g}^{\top}\mathbf{f}_{g}(t)=A_{g}\sqrt{4\pi^{2}+\gamma_{g}^{2}}\sin(2\pi t+\theta_{g}+\phi_{g})+b_{g}\gamma_{g},\quad\phi_{g}=\arctan(2\pi/\gamma_{g}).$
Parameters are optimized by matching predicted trajectories $\mathbf{\hat{x}}_{c}=(\mathbf{w}_{g}^{\top}\mathbf{f}_{g}(t_{g,c}),s_{g}(t_{g,c}))$ to observed data $\mathbf{x}_{c}=(\mathbf{w}_{g}^{\top}\mathbf{f}_{g,c},s_{g,c})$ using the same likelihood framework as Eq. 3.5.
Deep-learning-based RNA Velocity
Recently, the application of deep learning methods has expanded the possibilities for RNA Velocity estimation (Oller-Moreno et al., 2021; Raimundo et al., 2021). In RNA velocity analysis, the expressive power of neural networks is especially useful in inferring the latent state of cells, as well as encouraging the consistency of the learned vector field.
@@
The Variational Autoencoder (VAE) (Kingma et al., 2013) is an effective approach to model the distribution of data through latent variables to achieve high-dimensional data reconstruction. Its core idea is to introduce a latent variable $\mathbf{z}$ and express the distribution of data $\mathbf{x}$ as the following conditional distribution $p(\mathbf{x})=\int p(\mathbf{x}|\mathbf{z})p(\mathbf{z})\mathrm{d}\mathbf{z},$ where the generative process is $\mathbf{z}\sim p(\mathbf{z}),\quad\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}).$ The VAE employs a decoder $p_{\theta}(\mathbf{x}|\mathbf{z})$ and an encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ to map between the latent space and data space. The training objective is the Evidence Lower Bound (ELBO)
@@
$\mathcal{L}_{\mathrm{ELBO}}=\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{reg}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]-\mathrm{D}_{\mathrm{KL}}[q_{\phi}(\mathbf{z}|\mathbf{x})\|p(\mathbf{z})],\quad\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x}),$
@@
where the first term is the reconstruction loss, enforcing similarity between decoded data and observations, and the second term acts as a regularizer, aligning the latent distribution with the prior. Typically, the decoder outputs the mean $\mathbf{\mu}_{\theta}(\mathbf{z})$ of a Gaussian distribution with fixed variance $\sigma^{2}$, leading to $\mathcal{L}_{\mathrm{rec}}\propto-\dfrac{1}{2\sigma^{2}}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\|\mathbf{x}-\mathbf{\mu}_{\theta}(\mathbf{z})\|^{2}\right].$ Several methods are based on VAE to improve the RNA velocity model by taking advantage of the latent space.
@@
VeloAE (Qiao and Huang, 2021) computes RNA velocity using latent space representations. Its encoder maps the spliced RNA matrix $\mathbf{S}\in\mathbb{R}^{n_{c}\times n_{g}}$ and unspliced RNA matrix $\mathbf{U}\in\mathbb{R}^{n_{c}\times n_{g}}$ to latent representations $\tilde{\mathbf{S}}\in\mathbb{R}^{n_{c}\times d_{z}}$ and $\tilde{\mathbf{U}}\in\mathbb{R}^{n_{c}\times d_{z}}$, while the decoder reconstructs $\hat{\mathbf{S}}$ and $\hat{\mathbf{U}}$. Velo AE enforces the steady-state constraint on latent representations $\mathbf{\tilde{u}}_{i},\mathbf{\tilde{s}}_{i}$, resulting in a composite loss:
$\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{reg}}=\sum_{i=1}^{d_{z}}\mathrm{MSE}(\mathbf{\tilde{u}}_{i}-\gamma_{i}\mathbf{\tilde{s}}_{i})+\left[\mathrm{MSE}(\mathbf{S},\mathbf{\hat{S}})+\mathrm{MSE}(\mathbf{U},\mathbf{\hat{U}})\right].$
The RNA velocity is derived as $\mathbf{\tilde{u_{i}}}-\gamma_{i}\mathbf{\tilde{s_{i}}}$ after training.
@@
LatentVelo (Farrell et al., 2023) incorporates latent variables $\mathbf{z}_{c}=(\mathbf{u}^{(z)}_{c},\mathbf{s}^{(z)}_{c})$ and pseudotime $t_{c}$, representing latent-space unspliced/spliced RNA levels and cellular pseudotime. It assumes the dynamics:
$\dfrac{\mathrm{d}\mathbf{u}^{(z)}_{c}(t)}{\mathrm{d}t}$ $=f_{u}(\mathbf{u}^{(z)}_{c}(t),\mathbf{r}^{(z)}_{c}(t)),$
$\dfrac{\mathrm{d}\mathbf{s}^{(z)}_{c}(t)}{\mathrm{d}t}$ $=f_{s}(\mathbf{u}^{(z)}_{c}(t),\mathbf{s}^{(z)}_{c}(t)),$
$\dfrac{\mathrm{d}\mathbf{r}^{(z)}_{c}(t)}{\mathrm{d}t}$ $=f_{r}(\mathbf{s}^{(z)}_{c}(t),\mathbf{r}^{(z)}_{c}(t),\mathbf{h}_{c}),$
$\mathbf{h}_{c}$ $=f_{h}(\mathbf{s}^{(z),\mathrm{obs}}_{c},\mathbf{u}^{(z),\mathrm{obs}}_{c}),$
@@
where $\mathbf{s}^{(z),\mathrm{obs}}_{c}$ and $\mathbf{u}^{(z),\mathrm{obs}}_{c}$ are latent-space observations, $\mathbf{r}^{(z)}_{c}$ governs chromatin dynamics, and $f_{u},f_{s},f_{r},f_{h}$ are neural networks with $f_{h}$ computing the cell state encoding $\mathbf{h}_{c}$. Beyond standard VAE losses, an evolution loss ensures correct dynamics:
@@
$\mathcal{L}_{\mathrm{evol}}=\sum_{c=1}^{n_{c}}\mathbb{E}_{t_{c}\sim q(t_{c}|\mathbf{x})}\left[\dfrac{\|\mathbf{z}^{\mathrm{obs}}_{c}-\mathbf{z}_{c}(t_{c})\|^{2}+\|\mathbf{x}_{c}-\mathbf{\hat{x}}_{c}(t_{c})\|}{\sigma^{2}}\right].$
@@
VeloVI (Gayoso et al., 2024) models gene-specific state distributions $\boldsymbol{\pi}_{g,c}$ with a Dirichlet prior $\boldsymbol{\pi}_{gc}\sim\mathrm{Dirichlet}(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$, where states $k_{g,c}\in\{1,2,3,4\}$ correspond to induction, repression, induction steady, and repression steady. Key parameters include gene-specific rates $\alpha_{g},\beta_{g},\gamma_{g}$, pseudotime $t_{g,c}$, and switching time $t_{g}^{s}$. For repression-related states, $\alpha_{g}=0$ is fixed. Genes in transient states follow Eq. 3.4, while steady states use analytic solutions. Reconstruction losses follow standard VAE training.
VeloVAE (Gu et al., 2022) uses latent variables $\mathbf{z}_{c}\sim\mathcal{N}(0,I)$ and pseudotime $t_{c}\sim\mathcal{N}(t_{0},\sigma_{0}^{2})$. A fully connected network maps $\mathbf{z}$ to gene-specific parameters $\alpha_{c,g},\beta_{c,g},\gamma_{c,g}$, enabling reconstruction via Eq. 3.4 and training with standard VAE losses.

### v2

@@ 3.1.3 Continuous Dynamics Modeling
RNA Velocity Model and Parameter Estimation Based on the unspliced RNA and spliced RNA counts $u_{g}$ and $s_{g}$ for each gene, an underlying ODE model could be naturally derived based on mass-action law such that
$\dfrac{\mathrm{d}u_{g}}{\mathrm{d}t}$ $=\alpha_{g}(t)-\beta_{g}(t)u_{g}(t),$ (4)
@@
where $\alpha_{g}(t)$, $\beta_{g}(t)$, and $\gamma_{g}(t)$ represent the rates of mRNA transcription, splicing, and degradation, respectively. Here, $v_{g}=\dfrac{\mathrm{d}s_{g}}{\mathrm{d}t}$ is defined as the RNA velocity of gene $g$ La Manno et al. (2018). By concatenating the RNA velocities of all genes in a cell, a vector $\boldsymbol{v}=(v_{1},v_{2},\cdots,v_{n})$ is formed, which contains information about how the amounts of spliced RNA in the cell are changing. This vector represents the potential direction of the cell state evolution and can be used for downstream tasks of cell fate inference.
From an algorithm perspective, the central issue in the RNA velocity is to determine the parameters of the Equation (4) from static snapshot data, where the time $t$ of each cell is not explicitly known. In the following sections, we will summarize methods for solving $v_{g}$ and downstream tasks that utilize $v_{g}$.
Steady-State Assumption: Parameter Estimation in Velocyto La Manno et al. (2018)
In the original RNA velocity paper La Manno et al. (2018), parameter estimation was performed using linear regression with steady-state assumption to avoid the reliance on latent time $t$. Firstly, it assumes that for all genes $g$, $\alpha_{g}(t)$, $\beta_{g}(t)$, and $\gamma_{g}(t)$ are time-invariant. Secondly, it assumes that all genes share the same splicing rate $\beta$. Denote $\tilde{\alpha}=\dfrac{\alpha}{\beta}$ and $\tilde{\gamma}=\dfrac{\gamma}{\beta}$. To compute RNA velocity, one only needs to estimate $\tilde{\gamma}_{g}$ with the steady-state assumption that $\frac{ds_{g}}{dt}=0$. Indeed, we have the linear relation $\tilde{\gamma}_{g}=\dfrac{u_{g}(t)}{s_{g}(t)},$ suggesting that under the steady-state assumption, $\tilde{\gamma}_{g}$ can be estimated using linear regression. In practice, since most cells do not satisfy this assumption, it is commonly assumed that cells in the upper-right or lower-left regions of a scatter plot with unspliced RNA on the x-axis and spliced RNA on the y-axis are in equilibrium. Therefore, the common algorithm implementation is to limit the linear regression to the top or bottom $5\%$ of cells based on unspliced and spliced RNA levels.
If the dynamic equations Equation (4) are expressed probabilistically, the parameter estimation could be enhanced based on the linear regression formulation of steady-state stochastic model Bergen et al. (2020). The equation could be expressed as the regression problem
$\begin{bmatrix}\langle u_{g}(t)\rangle\\ \langle u_{g}(t)\rangle+2\langle u_{g}(t)s_{g}(t)\rangle\end{bmatrix}=\tilde{\gamma}_{g}\begin{bmatrix}\langle s_{g}(t)\rangle\\ 2\langle s_{g}^{2}(t)\rangle-\langle s_{g}(t)\rangle\end{bmatrix}+\boldsymbol{\epsilon},$
@@
Dynamic Inference: Parameter Estimation in scVelo Bergen et al. (2020)
A major issue with steady-state analysis is that many transient cells would be discarded in the parameter estimation. The scVelo approach Bergen et al. (2020) circumvents the issues through the estimation of kinetic parameters ($\alpha_{g}$, $\beta_{g}$, $\gamma_{g}$) via an expectation-maximization (EM) algorithm by modeling all the cells with the dynamical process.
Guided by transcriptional regulation principles, scVelo models gene expression dynamics through two distinct transcriptional phases: (1) an induction phase ($k=0$) characterized by promoter activation and transcriptional upregulation, and (2) a subsequent repression phase ($k=1$) marked by transcriptional suppression. This phase-specific regulation manifests through different unspliced RNA production rates ($\alpha^{(0)}_{g}\neq\alpha^{(1)}_{g}$). Let $t^{(k)}_{g}$ denote the transition time from phase $k-1$ to $k$ for gene $g$, with initial conditions $u_{g}^{\star}=u_{g}(t^{(k)}_{g})$ and $s_{g}^{\star}=s_{g}(t^{(k)}_{g})$. The analytical solution to Equation (4) during phase $k$ yields
@@
For each gene $g$, one can estimate the parameter set $\boldsymbol{\theta}_{g}=\{\alpha^{(k)}_{g},\beta_{g},\gamma_{g},t^{(k)}_{g}\}$ by minimizing the discrepancy between modeled trajectories $\boldsymbol{\hat{x}}_{g}(t)=(u_{g}(t),s_{g}(t))$ and observed single-cell measurements $\boldsymbol{x}_{g,c}=(u_{g,c},s_{g,c})$ across cells $c$. Assuming Gaussian residuals $\boldsymbol{e}_{c}=\|\boldsymbol{x}_{g,c}-\boldsymbol{\hat{x}}_{g}(t_{c})\|$ with variance $\sigma^{2}$, the log-likelihood function becomes
$\max_{\boldsymbol{\theta_{g}},t_{c}}\mathcal{L}(\boldsymbol{\theta}_{g},,t_{c})=-\frac{1}{2\sigma^{2}}\sum_{c}\|\boldsymbol{x}_{g,c}-\boldsymbol{\hat{x}}_{g}(t_{c})\|^{2}+\text{constant}.$ (5)
@@
    $\beta_{g}$ $=1,\quad\gamma_{g}=\frac{\boldsymbol{u}_{g}^{\top}\boldsymbol{s}_{g}}{\|\boldsymbol{s}_{g}\|^{2}},$
@@
  • E-step: Assigning hidden latent time $t_{c}$ for each cell by projecting observations onto the current estimated trajectory $\boldsymbol{\hat{x}}_{g}(t|\boldsymbol{\theta}_{g})$.
  • M-step: Updating $\boldsymbol{\theta}_{g}$ via maximum likelihood estimation given current latent time assignments.
@@
While traditional RNA velocity methods focus on estimating parameters $\alpha_{g},\beta_{g},\gamma_{g}$ in dynamic Equation (4), alternative approaches such as UniTVelo Gao et al. (2022) and TF Velo Li et al. (2024) take a different path by directly parameterizing the dynamics of spliced RNA. UniTVelo Gao et al. (2022) models transcriptional phases through radial basis functions
@@
where the velocity derives directly from function differentiation $\dot{s}_{g}(t_{g,c})=-2a_{g}(t_{g,c}-\tau_{g})s_{g}(t_{g,c}).$ The full parameter set $(h_{g},a_{g},\tau_{g},o_{g},\gamma_{g},\beta_{g},i_{g},t_{g,c})$ is estimated via maximum likelihood framework as Equation (5), comparing model predictions $\boldsymbol{\hat{x}}_{c}=(u_{g}(t_{g,c}),s_{g}(t_{g,c}))$ against observations $\boldsymbol{x}_{c}=(u_{g,c},s_{g,c})$ under Gaussian residuals.
TF Velo Li et al. (2024) introduces transcription factor coupling through linear dynamics $\dot{s}_{g}(t)=\boldsymbol{w}_{g}^{\top}\boldsymbol{f}_{g}(t)-\gamma_{g}s_{g}(t),$ combined with sinusoidal splicing $s_{g}(t)=A_{g}\sin(\omega_{g}t+\theta_{g})+b_{g}.$ This functional specification enables the analytical resolution of TF interactions
$\boldsymbol{w}_{g}^{\top}\boldsymbol{f}_{g}(t)=A_{g}\sqrt{4\pi^{2}+\gamma_{g}^{2}}\sin(2\pi t+\theta_{g}+\phi_{g})+b_{g}\gamma_{g},\quad\phi_{g}=\arctan(2\pi/\gamma_{g}).$
Parameters are optimized by matching predicted trajectories $\boldsymbol{\hat{x}}_{c}=(\boldsymbol{w}_{g}^{\top}\boldsymbol{f}_{g}(t_{g,c}),s_{g}(t_{g,c}))$ to observed data $\boldsymbol{x}_{c}=(\boldsymbol{w}_{g}^{\top}\boldsymbol{f}_{g,c},s_{g,c})$ using the same likelihood framework as Equation (5).
Deep Learning-Based RNA Velocity Recently, the application of deep learning methods has expanded the possibilities for RNA velocity estimation Oller-Moreno et al. (2021); Raimundo et al. (2021); Gu et al. (2022). In RNA velocity analysis, the expressive power of neural networks is especially useful in inferring the latent state of cells, as well as encouraging the consistency of the learned vector field.
@@
The Variational Autoencoder (VAE) Kingma et al. (2013) is an effective approach to model the distribution of data through latent variables to achieve high-dimensional data reconstruction. Its core idea is to introduce a latent variable $\boldsymbol{z}$ and express the distribution of data $\boldsymbol{x}$ as the following conditional distribution: $p(\boldsymbol{x})=\int p(\boldsymbol{x}|\boldsymbol{z})p(\boldsymbol{z})\mathrm{d}\boldsymbol{z},$ where the generative process is $\boldsymbol{z}\sim p(\boldsymbol{z}),\quad\boldsymbol{x}\sim p(\boldsymbol{x}|\boldsymbol{z}).$ The VAE employs a decoder $p_{\theta}(\boldsymbol{x}|\boldsymbol{z})$ and an encoder $q_{\phi}(\boldsymbol{z}|\boldsymbol{x})$ to map between the latent space and data space. The training objective is the Evidence Lower Bound (ELBO):
@@
$\mathcal{L}_{\mathrm{ELBO}}=\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{reg}}=\mathbb{E}_{q_{\phi}(\boldsymbol{z}|\boldsymbol{x})}[\log p_{\theta}(\boldsymbol{x}|\boldsymbol{z})]-\mathrm{D}_{\mathrm{KL}}[q_{\phi}(\boldsymbol{z}|\boldsymbol{x})\|p(\boldsymbol{z})],\quad\boldsymbol{x}\sim p_{\mathrm{data}}(\boldsymbol{x}),$
@@
where the first term is the reconstruction loss, enforcing similarity between decoded data and observations, and the second term acts as a regularizer, aligning the latent distribution with the prior. Typically, the decoder outputs the mean $\boldsymbol{\mu}_{\theta}(\boldsymbol{z})$ of a Gaussian distribution with fixed variance $\sigma^{2}$, leading to $\mathcal{L}_{\mathrm{rec}}\propto-\dfrac{1}{2\sigma^{2}}\mathbb{E}_{q_{\phi}(\boldsymbol{z}|\boldsymbol{x})}\left[\|\boldsymbol{x}-\boldsymbol{\mu}_{\theta}(\boldsymbol{z})\|^{2}\right].$ Several methods are based on VAE to improve the RNA velocity model by taking advantage of the latent space.
@@
VeloAE Qiao and Huang (2021) computes RNA velocity using latent space representations. Its encoder maps the spliced RNA matrix $\boldsymbol{S}\in\mathbb{R}^{n_{c}\times n_{g}}$ and unspliced RNA matrix $\boldsymbol{U}\in\mathbb{R}^{n_{c}\times n_{g}}$ to latent representations $\tilde{\boldsymbol{S}}\in\mathbb{R}^{n_{c}\times d_{z}}$ and $\tilde{\boldsymbol{U}}\in\mathbb{R}^{n_{c}\times d_{z}}$, while the decoder reconstructs $\hat{\boldsymbol{S}}$ and $\hat{\boldsymbol{U}}$. Velo AE enforces the steady-state constraint on latent representations $\boldsymbol{\tilde{u}}_{i},\boldsymbol{\tilde{s}}_{i}$, resulting in a composite loss:
$\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{reg}}=\sum_{i=1}^{d_{z}}\mathrm{MSE}(\boldsymbol{\tilde{u}}_{i}-\gamma_{i}\boldsymbol{\tilde{s}}_{i})+\left[\mathrm{MSE}(\boldsymbol{S},\boldsymbol{\hat{S}})+\mathrm{MSE}(\boldsymbol{U},\boldsymbol{\hat{U}})\right].$
The RNA velocity is derived as $\boldsymbol{\tilde{u_{i}}}-\gamma_{i}\boldsymbol{\tilde{s_{i}}}$ after training.
@@
LatentVelo Farrell et al. (2023) incorporates latent variables $\boldsymbol{z}_{c}=(\boldsymbol{u}^{(z)}_{c},\boldsymbol{s}^{(z)}_{c})$ and pseudotime $t_{c}$, representing latent-space unspliced/spliced RNA levels and cellular pseudotime. It assumes the following dynamics:
$\dfrac{\mathrm{d}\boldsymbol{u}^{(z)}_{c}(t)}{\mathrm{d}t}$ $=f_{u}(\boldsymbol{u}^{(z)}_{c}(t),\boldsymbol{r}^{(z)}_{c}(t)),$
$\dfrac{\mathrm{d}\boldsymbol{s}^{(z)}_{c}(t)}{\mathrm{d}t}$ $=f_{s}(\boldsymbol{u}^{(z)}_{c}(t),\boldsymbol{s}^{(z)}_{c}(t)),$
$\dfrac{\mathrm{d}\boldsymbol{r}^{(z)}_{c}(t)}{\mathrm{d}t}$ $=f_{r}(\boldsymbol{s}^{(z)}_{c}(t),\boldsymbol{r}^{(z)}_{c}(t),\boldsymbol{h}_{c}),$
$\boldsymbol{h}_{c}$ $=f_{h}(\boldsymbol{s}^{(z),\mathrm{obs}}_{c},\boldsymbol{u}^{(z),\mathrm{obs}}_{c}),$
@@
where $\boldsymbol{s}^{(z),\mathrm{obs}}_{c}$ and $\boldsymbol{u}^{(z),\mathrm{obs}}_{c}$ are latent-space observations, $\boldsymbol{r}^{(z)}_{c}$ governs chromatin dynamics, and $f_{u},f_{s},f_{r},f_{h}$ are neural networks with $f_{h}$ computing the cell state encoding $\boldsymbol{h}_{c}$. Beyond standard VAE losses, an evolution loss ensures correct dynamics:
@@
$\mathcal{L}_{\mathrm{evol}}=\sum_{c=1}^{n_{c}}\mathbb{E}_{t_{c}\sim q(t_{c}|\boldsymbol{x})}\left[\dfrac{\|\boldsymbol{z}^{\mathrm{obs}}_{c}-\boldsymbol{z}_{c}(t_{c})\|^{2}+\|\boldsymbol{x}_{c}-\boldsymbol{\hat{x}}_{c}(t_{c})\|}{\sigma^{2}}\right].$
@@
VeloVI Gayoso et al. (2024) models gene-specific state distributions $\boldsymbol{\pi}_{g,c}$ with a Dirichlet prior $\boldsymbol{\pi}_{gc}\sim\mathrm{Dirichlet}(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$, where states $k_{g,c}\in\{1,2,3,4\}$ correspond to induction, repression, induction steady, and repression steady. Key parameters include gene-specific rates $\alpha_{g},\beta_{g},\gamma_{g}$, pseudotime $t_{g,c}$, and switching time $t_{g}^{s}$. For repression-related states, $\alpha_{g}=0$ is fixed. Genes in transient states follow Equation (4), while steady states use analytic solutions. Reconstruction losses follow standard VAE training.
VeloVAE Gu et al. (2022) uses latent variables $\boldsymbol{z}_{c}\sim\mathcal{N}(0,I)$ and pseudotime $t_{c}\sim\mathcal{N}(t_{0},\sigma_{0}^{2})$. A fully connected network maps $\boldsymbol{z}$ to gene-specific parameters $\alpha_{c,g},\beta_{c,g},\gamma_{c,g}$, enabling reconstruction via Equation (4) and training with standard VAE losses.